# Graph Neural Network-based structural classification of glass-forming liquids and its interpretation via Self-Attention mechanism

### v1

@@ I Introduction
Specifically, ML methodologies have been applied for data obtained from molecular dynamics (MD) simulations to explore multifaceted origins of the slow dynamics from the atomistic level. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] These investigations have effectively opened new vistas for comprehending the characteristics between liquids and glasses, spanning the prediction of their properties to the elucidation of structural and dynamical changes by varying temperatures. The seamless integration of ML and MD represents a transformative approach, which yields insights into the behaviors inherent to various glass-forming liquids.
@@
More recently, there has been a notable increase in the application of sophisticated ML techniques, including Convolutional Neural Networks (CNNs) [44] and Graph Neural Networks (GNNs) [45, 46], as evidenced in literatures. Bapst et al. achieved remarkable success in accurately predicting the dynamic propensity in a three-dimensional glass-forming liquids from its static structure, utilizing the power of GNNs. [12] The dynamic propensity refers to the identification of single-particle displacements exhibiting higher mobility than the average. [47, 48, 49] Consequently, their work established a compelling link between the dynamics and structural attributes of glass-forming liquids. Shiba et al. also succeeded in predicting the dynamic propensity with higher accuracy than the results of Bapst et al. [25] Their success can be attributed to an extended approach that involved learning not only individual particle displacements but also the pair-particle displacements as part of the GNNs output. Additionally, other advanced GNN techniques were described, including the incorporation of a geometry-enhanced GNN [33] and a rotation-equivariant graph structure. [35]
@@
Swanson et al. pursued an alternative approach and found that CNNs and GNNs possess the capability to discriminate between liquid and glass structures across the glass transition temperature $T_{\mathrm{g}}$ of the two-dimensional glass-forming liquids. [13] Notably, the Self-Attention mechanism, a method within eXplainable Artificial Intelligence (XAI), provides insights into how GNNs differentiate liquids and glasses. Originally introduced in Graph Attention Networks, the Self-Attention mechanism quantifies the importance of a given node in relation to its edge-connected neighbors within the graph structure. [50, 51] Other studies have also highlighted the effectiveness of the Self-Attention mechanism in ML, revealing connections between structure features and dynamic heterogeneities. [33, 36, 41]
@@ II.2 GNN and Self-Attention mechanism
In general, the GNN is a neural network model designed to exploit the capabilities of ML within graph-structured data. It exploits the inherent nature of graph data as its input and iteratively updates information associated with the nodes and edges present in the input graph through message passing. Therefore, the GNN eliminates the requirement for introducing features manually, allowing automatic discovery of features. This GNN approach becomes particularly appropriate when applied to many particle systems since it effectively retains the three-dimensional structural information.
The schematic picture of our GNN architecture is depicted in Fig. 1. The particle configuration can be represented by embedding features of particle types and connections between pairs of particles as attributes of nodes and edges of the graph structure, respectively. Specifically, we utilize the particle type (type 1 or type 2) as input feature vectors for the graph nodes. A pair of particles are treated as neighbors, forming an edge when their distance is smaller than $r_{\mathrm{cut}}=1.8$. The relative coordinates of the pair of neighboring particles serve as input features for the edges. The value of $r_{\mathrm{cut}}=1.8$ corresponds to the first minimum length of the radial distribution functions $g(r)$ of large particles (type 2).
@@
In contrast to the approaches utilized in prior GNN studies for learning the dynamic propensity, we introduce the Self-Attention layer into our GNN architecture in order to assess which node’s information to emphasize, considering the information exchanged among neighboring nodes. [50, 51] Note that Swanson et al. employed the Self-Attention layer at the end of message passing in Ref. 13, which is compatible with our GNN architecture, as shown in Fig. 1. The attention coefficient $e_{vu}$ is defined as
@@ II.3 Structural descriptors for local particle environment
Note that the coarse-grained BOO has been proposed by averaging $q_{lm}(i)$ over neighboring particles of particle $i$ and particle $i$ itself. [65] This approach has been shown to provide a more precise and discriminative characterization between liquid and crystal structures when compared to the original BOO method. We have evaluated the use of this coarse-grained BOO in our neural network; however, we have opted not to employ it in this study. Our observations indicated that the additional averaging step tends to obscure the distinctions among disordered structural configurations within the systems we are investigating. An analogous observation has been documented in a previous study, where unsupervised machine learning was employed for dimension reduction of local structural orders in glass-forming liquids. [22] An alternative approach to characterize local orders is based on Voronoi tessellation. [60] In particular, the Voronoi cell serves as an indicator of the local geometry associated with each particle in glass-forming liquids. [66, 67] We conducted the Voronoi cell volume calculations using the Pyvoro code. [68] A final structural descriptor examined in this study is C.N. within the first coordination shell, defined as the number of particles within $r_{\mathrm{cut}}=1.8$.
@@ III.2 Node features generated by GNN
We next examined the physical properties of the node features generated by the GNN, denoted as $H^{\prime}_{u}$ for the node index $u$ in Fig. 1, which were were subsequently employed in the FCNN to predict temperature. For this aim, we calculated BOOs ($Q_{6}$ and $Q_{4}$), Voronoi cell colume, and C.N. for each particle of a specific configuration and compared these structural descriptors with the GNN-generated node features $H^{\prime}_{u}$ using the same configuration. Specifically, the Pearson correlation coefficient was calculated for each $(T_{1},T_{2})$ pair snapshot and averaged over 200 configurations (100 configurations from $T_{1}$ and 100 configurations of $T_{2}$ from the test dataset). The results are illustrated in Fig. 3. Overall, $Q_{6}$ and Voronoi cell volumes display positive correlations with the node features, while C.N. exhibit negative correlations. Note that $Q_{4}$ also exhibits positive correlations, albeit less prominently than $Q_{6}$. A more detailed observation reveals that, for all ($T_{1}$, $T_{2}$) pairs, the absolute values of the correlations with node features generated by GNNs follow the order: C.N. $>$ Voronoi cell volume $>$ $Q_{6}$ $>$ $Q_{4}$. This trend can be attributed to the methodology used to compute structural descriptors: $Q_{6}$ and $Q_{4}$ incorporate angular information of neighboring particles. In contrast, C.N. and Voronoi cell volume provide a more direct capture of local structural environment. The observed correlations between node features and structural descriptors suggest that the GNN’s success in two-temperature discrimination arises from its ability to effectively capture insights into the local particles environment.
@@ III.3 FCNN classification of different temperatures using structural descriptors
When $Q_{6}$ is used as the input variable, the structure at the lowest temperature, $T_{1}=0.21$, is clearly distinguished from those at other temperatures. However, distinguishing structures between high temperatures proves challenging, in contrast to the GNN results. When other structural descriptors, $Q_{4}$, Voronoi cell volume, and C.N., are used as input variables, the predictive performance follows a similar trend to that of $Q_{6}$, albeit with an overall degradation. These reductions in performance, particularly in the high-temperature regime, suggest that using physically-defined structural descriptors as input variables in the FCNN presents certain limitations. The comparison between the GNN and FCNN results across the four types for structural descriptors as input variables highlights the strength of the GNN approach, since it operates independently of prior knowledge regarding the structural properties of the systems under investigation.
@@ IV Conclusions
The adoption of GNNs proved highly advantageous, given their innate capability to process graph data efficiently. By representing individual particles as nodes and their connections as edges within our graph-based framework, we utilized the advantages of GNNs to automatically generate feature representations, without relying on prior physical information. These features captured intricate relationships with neighboring particles and allowed us to discern distinct temperature-dependent structures. Notably, node features generated by the GNN exhibit correlations with physically-defined structural descriptors including the BOOs, $Q_{6}$ and $Q_{4}$, Voronoi cell volume, and C.N. The results demonstrates that the GNN learns information related to the local structure of the particle configuration from the amorphous structure.
Furthermore, we conducted a comparative analysis by applying conventional FCNNs to the same binary classification problem. Using $Q_{6}$, $Q_{4}$, Voronoi cell volume, and C.N. as input variables, this approach aimed to provide deeper insights into the fundamental mechanisms underpinning ML in glass-forming liquids. This analysis not only corroborated the significance of structural descriptors but also highlighted the distinct strengths of GNNs in capturing important features from amorphous configurations without requiring explicit prior knowledge.
@@
Our exploration of GNNs for structural classification in glass-forming liquids opens several promising aspects for future research. Extending this work to encompass a broader range of materials, including different glass formers such as the widely used Kob–Andersen model, silica glasses, polymer glasses and supercooled water, will contribute to a more comprehensive understanding of the relationships between structural features and dynamics. In particular, the integration of additional structural descriptors parameters [70, 71], in conjunction with the GNN+Self-Attention framework, will provide valuable insights into the underlying mechanisms in glass-forming liquids.

### v2

@@ I Introduction
Specifically, ML methodologies have been applied for data obtained from molecular dynamics (MD) simulations to explore multifaceted origins of the slow dynamics from the atomistic level. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] These investigations have effectively opened new perspectives for comprehending the characteristics between liquids and glasses, spanning the prediction of their properties to the elucidation of structural and dynamical changes by varying temperatures. The seamless integration of ML and MD represents a transformative approach, which yields insights into the behaviors inherent to various glass-forming liquids.
@@
More recently, there has been a notable increase in the application of sophisticated ML techniques, including Convolutional Neural Networks (CNNs) [44] and Graph Neural Networks (GNNs) [45, 46], as evidenced in literatures. In particular, the GNN is a neural network model designed to exploit the capabilities of ML within graph structure, where the information is represented in terms of nodes and edges. It iteratively updates information associated with the nodes and edges present in the input graph through message passing. Therefore, the GNN eliminates the requirement for introducing features manually, allowing automatic discovery of relevant structural information. Bapst et al. achieved remarkable success in accurately predicting the dynamic propensity in a three-dimensional glass-forming liquids from its static structure, utilizing the power of GNNs. [12] The dynamic propensity refers to the identification of single-particle displacements exhibiting higher mobility than the average. [47, 48, 49] Consequently, their work established a compelling link between the dynamics and structural attributes of glass-forming liquids. Shiba et al. also succeeded in predicting the dynamic propensity with higher accuracy than the results of Bapst et al. [25] Their success can be attributed to an extended approach that involved learning not only individual particle displacements but also the pair-particle displacements as part of the GNNs output. Additionally, other advanced GNN techniques were described, including the incorporation of a geometry-enhanced GNN [33] and a rotation-equivariant graph structure. [35] These studies demonstrated that the GNNs can automatically extract features from graph-based inputs through message passing, without the need for predefined physical descriptors. However, their interpretability remains limited due to their black-box nature.
@@
Swanson et al. pursued an alternative approach and found that CNNs and GNNs possess the capability to discriminate between liquid and glass structures across the glass transition temperature $T_{\mathrm{g}}$ of the two-dimensional glass-forming liquids. [13] Notably, the self-attention mechanism, a method within eXplainable Artificial Intelligence (XAI), provides insights into how GNNs differentiate liquids and glasses. Originally introduced in Graph Attention Networks, the self-attention mechanism quantifies the importance of a given node in relation to its edge-connected neighbors within the graph structure, providing interpretability to black-box nature of GNN. [50, 51] Other studies have also highlighted the effectiveness of the self-attention mechanism in ML, revealing connections between structure features and dynamic heterogeneities. [33, 36, 41]
@@ II.2 GNN and self-attention mechanism
The schematic picture of our GNN architecture is depicted in Fig. 1. The particle configuration can be represented by embedding features of particle types and connections between pairs of particles as attributes of nodes and edges of the graph structure, respectively. Specifically, we utilize the particle type (type 1 or type 2) as input feature vectors for the graph nodes. A pair of particles are treated as neighbors, forming an edge when their distance is smaller than $r_{\mathrm{cut}}=1.8$. The relative coordinates of the pair of neighboring particles serve as input features for the edges. The value of $r_{\mathrm{cut}}=1.8$ corresponds to the first minimum length of the radial distribution functions $g(r)$ of large particles (type 2). Note that this GNN approach is particularly appropriate when applied to many particle systems since it effectively retains the three-dimensional structural information.
@@
We introduce the self-attention layer into our GNN architecture in order to assess which node’s information to emphasize, considering the information exchanged among neighboring nodes. [50, 51] Note that Swanson et al. employed the self-attention layer at the end of message passing in Ref. 13, which is compatible with our GNN architecture, as shown in Fig. 1. The attention coefficient $e_{vu}$ is defined as
@@ II.3 Structural descriptors for local particle environment
Note that the coarse-grained BOO has been proposed by averaging $q_{lm}(i)$ over neighboring particles of particle $i$ and particle $i$ itself. [65] This approach has been shown to provide a more precise and discriminative characterization between liquid and crystal structures when compared to the original BOO method. We have evaluated the use of this coarse-grained BOO in our neural network; however, we have opted not to employ it in this study. Our observations indicated that the additional averaging step tends to obscure the distinctions among disordered structural configurations within the systems we are investigating. Note that an analogous observation has been documented in a previous study, where unsupervised machine learning was employed for dimension reduction of local structural orders in glass-forming liquids. [22]
An alternative approach to characterize local orders is based on Voronoi tessellation. [60] In particular, the Voronoi cell serves as an indicator of the local geometry associated with each particle in glass-forming liquids. [66, 67] We conducted the Voronoi cell volume calculations using the Pyvoro code. [68] A final structural descriptor examined in this study is C.N. within the first coordination shell, defined as the number of particles within $r_{\mathrm{cut}}=1.8$.
@@ III.2 Node features generated by GNN
We next examined the physical properties of the node features generated by the GNN, denoted as $H^{\prime}_{u}$ for the node index $u$ in Fig. 1, which were were subsequently employed in the FCNN to predict temperature. For this aim, we calculated BOOs ($Q_{6}$ and $Q_{4}$), Voronoi cell colume, and C.N. for each particle of a specific configuration and compared these structural descriptors with the GNN-generated node features $H^{\prime}_{u}$ using the same configuration. Specifically, the Pearson correlation coefficient was calculated for each $(T_{1},T_{2})$ pair snapshot and averaged over 200 configurations (100 configurations from $T_{1}$ and 100 configurations of $T_{2}$ from the test dataset). The results are illustrated in Fig. 3. Overall, $Q_{6}$ and Voronoi cell volumes display positive correlations with the node features reaching values approximately 0.3, while C.N. exhibit negative correlations of up to 0.4 in magnitude. Note that $Q_{4}$ also shows slight positive correlations, though less procounced than those of $Q_{6}$. A more detailed observation reveals that, for all ($T_{1}$, $T_{2}$) pairs, the absolute values of the correlations with node features generated by GNNs follow the order: C.N. $>$ Voronoi cell volume $>$ $Q_{6}$ $>$ $Q_{4}$. This trend can be attributed to the methodology used to compute structural descriptors: $Q_{6}$ and $Q_{4}$ incorporate angular information of neighboring particles. In contrast, C.N. and Voronoi cell volume provide a more direct capture of local structural environment. The observed correlations between node features and structural descriptors suggest that the GNN’s success in two-temperature discrimination arises from its ability to effectively capture insights into the local particles environment.
@@ III.3 FCNN classification of different temperatures using structural descriptors
When $Q_{6}$ is used as the input variable, the structure at the lowest temperature, $T_{1}=0.21$, is clearly distinguished from those at other temperatures. However, distinguishing structures between high temperatures proves challenging, in contrast to the GNN results. When other structural descriptors, $Q_{4}$, Voronoi cell volume, and C.N., are used as input variables, the predictive performance follows a similar trend to that of $Q_{6}$, albeit with an overall degradation. These reductions in performance, particularly in the high-temperature regime, suggest that using physically-defined structural descriptors as input variables in the FCNN presents certain limitations. In particular, the results suggest that FCNNs with structural descriptors are more sensitive to thermal fluctuations at high temperatures. In contrast, the decrease in prediction accuracy for GNNs is relatively modest, implying that the GNN effectively suppresses the influence of thermal fluctuations and generates more robust features through the message passing process.
@@ IV Conclusions
The adoption of GNNs proved highly advantageous, given their inherent capability to process graph data efficiently. By representing individual particles as nodes and their connections as edges within our graph-based framework, we utilized the advantages of GNNs to automatically generate feature representations, without relying on prior physical information. These features captured intricate relationships with neighboring particles and allowed us to discern distinct temperature-dependent structures. Notably, node features generated by the GNN exhibit correlations with physically-defined structural descriptors including the BOOs, $Q_{6}$ and $Q_{4}$, Voronoi cell volume, and C.N. The results demonstrates that the GNN learns information related to the local structure of the particle configuration from the amorphous structure.
Furthermore, we conducted a comparative analysis by applying conventional FCNNs to the same binary classification problem. Using $Q_{6}$, $Q_{4}$, Voronoi cell volume, and C.N. as input variables, this approach aimed to provide deeper insights into the fundamental mechanisms underpinning ML in glass-forming liquids. This analysis not only corroborated the significance of structural descriptors but also highlighted the distinct strengths of GNNs in capturing important features from amorphous configurations.
@@
Our exploration of GNNs for structural classification in glass-forming liquids opens several promising aspects for future research. Extending this work to encompass a broader range of materials, including different glass formers such as the widely used Kob–Andersen model, silica glasses, polymer glasses and supercooled water, will contribute to a more comprehensive understanding of the relationships between structural features and dynamics. In particular, the integration of additional structural descriptors parameters [70, 71], in conjunction with the GNN+self-attention framework, will valuable insights into the underlying mechanisms in glass- forming liquids. For silica glasses and supercooled water, combining this approach with structural descriptors, which captures the degree of local tetrahedral ordering, may provide deeper insights into the structural origins of their anomalous behaviors. [72] In addition, incorporating persistent homology could further advance the analysis from a topological perspective, even in polymer glasses lacking a well-defined structural motif. [73]
SUPPLEMENTARY MATERIAL
The supplementary material includes temperature dependence of $g(r)$, $F_{\mathrm{s}}(k,t)$, $\tau_{\alpha}$, and potential energy (Fig. S1); probability distribution of structural descriptors, including the BOOs, Voronoi cell volume, and C.N. (Fig. S2); epoch-dependent loss function of GNNs (Fig. S3); fraction of GNN accuracy and AUC for the temperature range 0.21-0.25 (Fig. S4); epoch-dependent loss function of GNNs for the temperature range 0.21-0.25 (Fig. S5); epoch-dependent loss function of FNCCs using structural descriptors (Fig. S6, S7, S8, and S9); fraction of FCNN AUC (Fig. S10); Probability distribution plots using structural descriptors for two temperature combination, shown alongsidethe coordination-weighted attention coefficient, $\tilde{\alpha}_{vu}$, for comparison (Fig. S11, S12, S13, and S14).
@@ Acknowledgements.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Data availability statement
The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.15151938. Further data that support the findings of this study are available from the corresponding author upon request.